OpenStudy (anonymous):

Anyone good with matrices?? I have a couple questions I couldn't figure out on that quiz.

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

One factor in flood safety along a levee is the area that will absorb water should the levee break. The coordinates that make up the boundary area are (0,0), (2.8, 2.1), and (4.3, 4.4). What is the area of the land that would absorb the water??

OpenStudy (anonymous):

There are answer options. a. 1.65 mi^2 b. 9.03 mi^2 c. 1.1 mi^2 d. 10.68 mi^2

OpenStudy (amistre64):

determinant of a matrix ...

OpenStudy (anonymous):

could you please explain the question more elaborately

OpenStudy (amistre64):

(de+af+cb) a b 1 a b c d 1 c d e f 1 e f -(ad+be+cf) A =1/2 [(de+af+cb)-(ad+be+cf)]

OpenStudy (amistre64):

(0,0) a b (2.8, 2.1) c d (4.3, 4.4) e f anything with an ab is zero since a and b are zero sooo A =1/2 [(de+af+cb)-(ad+be+cf)] A =1/2 [(de)-(cf)]

OpenStudy (amistre64):

http://www.mathplanet.com/education/algebra-2/matrices/determinants using determinant to figure area of a traingle

OpenStudy (amistre64):

since we can move any triangle so that one of its verts are at the origin, this should amount to 1/2 the determinant of the other 2 points after adustments

OpenStudy (anonymous):

so A=1/2 (9.03 - 12.32) ?

OpenStudy (anonymous):

wait never mind that doesn't make sense, sorry I messed up

OpenStudy (amistre64):

yes, but its absolute value, so the results can ignore the sign

OpenStudy (anonymous):

oh.okay. I got it thank you so much!

OpenStudy (amistre64):

(2.8, 2.1) (4.3, 4.4) 1/2 (2.8(4.4) - 2.1(4.3)) ^^^^^^^^^^^^^ det of 2 nonzero points

OpenStudy (amistre64):

youre welcome